OpenStudy (anonymous):

Please help in solving : (2x+10)/((x)*(x+10))= 1/12 , find x

OpenStudy (anonymous):

12(2x+10)=x^2+10x x^2-14x-120=0 (x-20)(x+6)=0 x=20.-6

OpenStudy (anonymous):

any simpler way to solve it

OpenStudy (anonymous):

(2x+10)/(x(x+10))=1/12 (x+x+10)/(x(x+10))=1/12 1/(x+10)+1/x =1/12=1/20+1/30 this gives u x=20; and if u write 1/12=1/4-1/6 =1/(-6+10)+1/(-6) this gives u x=-6

OpenStudy (anonymous):

cold you please let me know how you got: 1/(x+10)+1/x =1/12=1/20+1/30

OpenStudy (anonymous):

1/30+1/20=(3+2)/60=5/60=1/12

OpenStudy (anonymous):

the best way to solve this is using factorization which i did first